There are 850 Douglas fir and ponderosa pine trees in a
section of forest bought by Sawz Logging Co. The company paid an average of
$300 for each Douglas fir and $225 for each ponderosa pine. If the company paid
$217,500 for the trees, how many of each kind did the company buy?

Answer by stanbon(57335)

There are 850 Douglas fir and ponderosa pine trees in a
section of forest bought by Sawz Logging Co. The company paid an average of
$300 for each Douglas fir and $225 for each ponderosa pine. If the company paid
$217,500 for the trees, how many of each kind did the company buy?
---------------
Let number of Douglas fir be "x"; Value of these is 300x dollars
Number of Ponderosa trees is "850-x"; Value of these is 225(850-x)=191250-250x
-----------
EQUATION:
value + value = 217,500
300x+191250-250=217,500
50x=26250
x=525 (# of Doublas fir trees)
850-525=325 (# of Ponderosa pine trees)
----------
Cheers,
Stan H.